OpenStudy (anonymous):

A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax and x represents the original price of the item. : C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale? Remember, do not include the $ symbol.

OpenStudy (anonymous):

C = 0.765x + 0.06(0.765x) C=.765*100+.0435*100 C=76.50+4.35 C=$80.85 IS THE DISCOUNTED PRICE.

OpenStudy (mathstudent55):

@Rachel_K Good job but check your arithmetic. 0.06 * 0.765 = 0.0459 You have 0.0435 I think you did 0.06 * 0.725 instead of 0.06 * 0.765 by mistake.

OpenStudy (anonymous):

thank y'all

OpenStudy (mathstudent55):

C = 0.765x + 0.06(0.765x) C = 0.765(100) + 0.0459(100) C = 76.50 + 4.59 C = 81.09

OpenStudy (anonymous):

oh yeah sry bout that part ... and yw @DeLinsko :)